OpenStudy (anonymous):

CHECK ANSWER are the power input and output in watts. If an amplifier has a power gain of 24 watts and the output power was 16 watts, find the power input in watts. in put power=.063

OpenStudy (anonymous):

input =16/24=.67watt

OpenStudy (anonymous):

output=input*gain

OpenStudy (anonymous):

is it the same formula as the earlier one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

which one are u talking of @him

OpenStudy (anonymous):

P(i) = 16/ (10^(2.4))